# for functionals in the $\mathcal{A}$-free setting

### v1

@@ 8. Stochastic homogenization
Let $\mathcal{F}^{\mathrm{st}}$ be the collection of functions $f\colon\Omega\times\mathbb{R}^{N}\times\mathbb{R}^{d}\to[0,+\infty)$ satisfying the following properties:
@@
    setting $f(\omega):=f(\omega,\cdot,\cdot)$, we have $f(\omega)\in\mathcal{F}$ for every $\omega\in\Omega$;
@@
We are now in a position to state our main result on stochastic homogenization of $\mathcal{A}$-free integral functionals. Given a stochastic integrand $f\in\mathcal{F}^{\mathrm{st}}$ and $\varepsilon>0$, we consider the stochastic integrands $f_{\varepsilon}\colon\Omega\times\mathbb{R}^{N}\times\mathbb{R}^{d}\to[0,+\infty)$ defined by
@@
for every $x\in\mathbb{R}^{N}$ and $\xi\in\mathbb{R}^{d}$. Setting $f_{\varepsilon}(\omega):=f_{\varepsilon}(\omega,\cdot,\cdot)$, we have that $f_{\varepsilon}(\omega)\in\mathcal{F}$ for every $\omega\in\Omega$, so we can consider the functionals $F_{\varepsilon}(\omega)\in\mathcal{I}$ associated with $f_{\varepsilon}(\omega)$ by (2.13), and the corresponding $\mathcal{A}$-free functionals $F_{\varepsilon}^{\mathcal{A}}(\omega)$ introduced in Definition 4.4.
@@
Let $f\in\mathcal{F}^{\mathrm{st}}$ and, for every $\varepsilon>0$ and for every $\omega\in\Omega$, let $F_{\varepsilon}(\omega)\in\mathcal{I}$ and $F_{\varepsilon}^{\mathcal{A}}(\omega)$ be the corresponding functionals. Then, there exists a set $\Omega^{\prime}\in{\mathcal{T}}$, with $P(\Omega^{\prime})=1$, such that for every $k\in\mathbb{N}$, $\omega\in\Omega^{\prime}$, and $\xi\in\mathbb{Q}^{d}$, the limit
@@
Then, $f_{\mathrm{hom}}\in\mathcal{F}^{\mathrm{st}}_{\mathrm{qc}}$ and, setting $f_{\mathrm{hom}}(\omega):=f_{\mathrm{hom}}(\omega,\cdot)$ for every $\omega\in\Omega$, the corresponding functionals $F_{\mathrm{hom}}(\omega)\in\mathcal{I}$ defined by (2.13) and $F_{\mathrm{hom}}^{\mathcal{A}}(\omega)$ introduced in Definition 4.4 satisfy the following properties:
@@
The existence of $\Omega^{\prime}\in{\mathcal{T}}$, with $P(\Omega^{\prime})=1$, such that (8.1) holds follows from Theorem 8.3 and Lemma 8.4. The properties of $f_{\mathrm{hom}}^{k}(\omega,\xi)$ and $f_{\mathrm{hom}}(\omega,\xi)$ for $\omega\in\Omega^{\prime}$ are given by Theorem 7.1, which implies also that (a) and (b) hold and that $f_{\mathrm{hom}}(\omega)\in\mathcal{F}_{\mathrm{qc}}$ for every $\omega\in\Omega$; hence, $f_{\mathrm{hom}}\in\mathcal{F}^{\mathrm{st}}_{\mathrm{qc}}$.
@@ Acknowledgements
The authors thank the hospitality of the Center of Nonlinear Analysis (CNA) at Carnegie Mellon University (Pittsburgh, USA) and of Erwin Schrödinger International Institute for Mathematics and Physics (Vienna, Austria).
The research of Gianni Dal Maso was partially supported by the National Research Project PRIN 2022J4FYNJ “Variational methods for stationary and evolution problems with singularities and interfaces” funded by the Italian Ministry of University and Research. Gianni Dal Maso is member of GNAMPA of INdAM.

### v2

@@ 8. Stochastic homogenization
Let $\mathcal{F}^{\mathrm{st}}_{\mathrm{Lip}}$ be the collection of functions $f\colon\Omega\times\mathbb{R}^{N}\times\mathbb{R}^{d}\to[0,+\infty)$ satisfying the following properties:
@@
    setting $f(\omega):=f(\omega,\cdot,\cdot)$, we have $f(\omega)\in\mathcal{F}_{\mathrm{Lip}}$ for every $\omega\in\Omega$;
@@
We recall that if the vector space $\mathrm{span}(\Lambda)$ generated by the wave cone $\Lambda$ coincides with $\mathbb{R}^{d}$ then $\mathcal{F}^{\mathrm{st}}_{\mathrm{qc}}\subset\mathcal{F}^{\mathrm{st}}_{\mathrm{Lip}}$ by Remark 2.8 .
@@
We are now in a position to state our main result on stochastic homogenization of $\mathcal{A}$-free integral functionals. Given a stochastic integrand $f\in\mathcal{F}^{\mathrm{st}}_{\mathrm{Lip}}$ and $\varepsilon>0$, we consider the stochastic integrands $f_{\varepsilon}\colon\Omega\times\mathbb{R}^{N}\times\mathbb{R}^{d}\to[0,+\infty)$ defined by
@@
for every $x\in\mathbb{R}^{N}$ and $\xi\in\mathbb{R}^{d}$. Setting $f_{\varepsilon}(\omega):=f_{\varepsilon}(\omega,\cdot,\cdot)$, we have that $f_{\varepsilon}(\omega)\in\mathcal{F}_{\mathrm{Lip}}$ for every $\omega\in\Omega$, so we can consider the functionals $F_{\varepsilon}(\omega)\in\mathcal{I}_{\mathrm{Lip}}$ associated with $f_{\varepsilon}(\omega)$ by (2.14), and the corresponding $\mathcal{A}$-free functionals $F_{\varepsilon}^{\mathcal{A}}(\omega)$ introduced in Definition 4.4.
@@
Let $f\in\mathcal{F}^{\mathrm{st}}_{\mathrm{Lip}}$ and, for every $\varepsilon>0$ and for every $\omega\in\Omega$, let $F_{\varepsilon}(\omega)\in\mathcal{I}_{\mathrm{Lip}}$ and $F_{\varepsilon}^{\mathcal{A}}(\omega)$ be the corresponding functionals. Then, there exists a set $\Omega^{\prime}\in{\mathcal{T}}$, with $P(\Omega^{\prime})=1$, such that for every $k\in\mathbb{N}$, $\omega\in\Omega^{\prime}$, and $\xi\in\mathbb{Q}^{d}$, the limit
@@
Then, $f_{\mathrm{hom}}\in\mathcal{F}^{\mathrm{st}}_{\mathrm{Lip}}\cap\mathcal{F}^{\mathrm{st}}_{\mathrm{qc}}$ and, setting $f_{\mathrm{hom}}(\omega):=f_{\mathrm{hom}}(\omega,\cdot)$ for every $\omega\in\Omega$, the corresponding functionals $F_{\mathrm{hom}}(\omega)\in\mathcal{I}_{\mathrm{Lip}}\cap\mathcal{I}_{\mathrm{qc}}$ defined by (2.14) and $F_{\mathrm{hom}}^{\mathcal{A}}(\omega)$ introduced in Definition 4.4 satisfy the following properties:
@@
The existence of $\Omega^{\prime}\in{\mathcal{T}}$, with $P(\Omega^{\prime})=1$, such that (8.1) holds follows from Theorem 8.3 and Lemma 8.4. The properties of $f_{\mathrm{hom}}^{k}(\omega,\xi)$ and $f_{\mathrm{hom}}(\omega,\xi)$ for $\omega\in\Omega^{\prime}$ are given by Theorem 7.1, which implies also that (a) and (b) hold and that $f_{\mathrm{hom}}(\omega)\in\mathcal{F}_{\mathrm{Lip}}\cap\mathcal{F}_{\mathrm{qc}}$ for every $\omega\in\Omega$; hence, $f_{\mathrm{hom}}\in\mathcal{F}^{\mathrm{st}}_{\mathrm{Lip}}\cap\mathcal{F}^{\mathrm{st}}_{\mathrm{qc}}$.
@@ Acknowledgements
The authors thank Jean-François Babadjian for the proof of Theorem 3.5, which provides the integral representation of $F$ without assuming a $p$-Lipschitz continuity. The authors thank the hospitality of the Center of Nonlinear Analysis (CNA) at Carnegie Mellon University (Pittsburgh, USA) and of Erwin Schrödinger International Institute for Mathematics and Physics (Vienna, Austria).
The research of Gianni Dal Maso was partially supported by the National Research Project PRIN 2022J4FYNJ “Variational methods for stationary and evolution problems with singularities and interfaces” funded by the Italian Ministry of University and Research. When this paper was written, Gianni Dal Maso was member of GNAMPA of INdAM.